TDAC has just paid $3 per share. The required rate of return for the stock is 9%. It had difficult few years and so the Board of Directors decided to cut the dividend by 3% every year, starting next dividend to be paid. What is the price of the stock at the scenario? There are some optimist investors too. They think the decrease will be for the next 2 years only and after that the dividend will again grow at constant growth rate of 5% thereafter. The required rate of return for the stock is again 9%. What is the price of the stock for optimist investor?
Answer #1

Current Dividend = D0 = $ 3, Growth Rate = - 3 %, Required Return = 9 %

Expected Dividend = 3 x (1-0.03) = $ 2.91

Current Stock Price = 2.91 / [0.09-(-0.03)] = $ 24.25

If the stock price falls for only 2 years and grows at 5% post that then dividends will be as shown below:

D1 = $ 2.91,D2 = 2.91 x 0.97 = $ 2.8227

D3 = 2.8227 x 1.05 = $ 2.963835

Current Stock Price = 2.91 / (1.09) + 2.8227 / (1.09)^(2) + [2.963835 / (0.09-0.05)] x [1/(1.09)^(2)] = $ 67.41